What is the force exerted Big Ben on the Empire State building? Assume that Big Ben has a mass of 10^8 kg and the Empire State building 10^9 kg. The distance between them is about 5000 kg and Big Ben is due east of the Empire State building.


1
Expert's answer
2021-12-20T10:27:03-0500

By the Newton's Gravitational law, we get:


"F=\dfrac{Gm_1m_2}{d^2},""F=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times10^8\ kg\times10^9\ kg}{(5\times10^6\ m)^2}=2.7\times10^{-7}\ N."